OpenStudy (anonymous):

ABCD is a parallelogram with a perimeter of 52. If ABNM is 36 and NC is congruent to AM. Find NM

OpenStudy (anonymous):



undeadknight26 (undeadknight26):

@helpme1.2 @jdoe0001

OpenStudy (anonymous):

AD = BC ----------------------- (1), AB = DC let AM = x = NC AB = DC = y BN = MD = p ..................... because of 1) ========================== 52 = 2x + 2y + 2p = 2[x+y+p] x +y+p = 26 ---------------------------------------- (2) ============================ y + p + x + NM = 36 26 + NM = 36 NM = 10

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

welcome